OpenStudy (anonymous):

Solve for x. 4(x-3)-11=15-2(x-4)

OpenStudy (anonymous):

First distribute the 4 and -2 to everything in the parentheses. So you'll get 4x-12-11=15-2x+8

OpenStudy (anonymous):

Combine like terms. 4x-23=-2x+23